OpenStudy (anonymous):

find the center and radius of the circle x^2+2x-4y+6=0 will FAN and MEDAL

OpenStudy (mathstudent55):

This is not a circle. Are you sure you copied the equation correctly? You need a y^2 term.

OpenStudy (anonymous):

@mathstudent55 That is literally whats on my test. The answer options are: center (-1,2) radius :11 center (-1,2) radius: sq root of 11 center (1,-2) radius: 11 center (1,-2) radius sq root of 11

OpenStudy (mertsj):

Well then let's work backwards from the answers and see which answer works.

OpenStudy (mertsj):

\[(x+1)^2+(y-2)^2=11^2\]

OpenStudy (mertsj):

\[x^2+2x+1+y^2-4y+4=121\]

OpenStudy (mertsj):

\[x^2+2x+y^2-4y=116\]

OpenStudy (mertsj):

Doesn't seem right, does it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I am confused I kind of get what you did but I don't know how to continue it

OpenStudy (mertsj):

The next choice has the same center but a different radius. But the radius should be sqrt116. So it doesn't seem to work either.

OpenStudy (mertsj):

So try the third choice.

OpenStudy (mertsj):

No. It couldn't be (x-1)^2 because that would be -2x

OpenStudy (mertsj):

So of the choices given it has to be the first or the second one. And we know it isn't the first one.

OpenStudy (mertsj):

Let's look at choice two more closely. I bet it will work.

OpenStudy (anonymous):

ok! after could you help me with another one? it's the same but the equation is x^2+8x-6y+9=0 options are center (4,-3) radius sq root of 34 center (-4,3) radius sq root of 34 center (4,-3) radius 34 center (-4,3) radius 34

OpenStudy (mertsj):

\[(x-1)^2+(y+2)^2=(\sqrt{11})^2\] \[x^2-2x+1+y^2+4y+4=11\] \[x^2-2x+y^2+4y+5=11\] \[x^2-2x+y^2+4y=6\]

OpenStudy (mertsj):

The question is obviously flawed if you have posted it correctly but since it is multiple choice, b is the best shot you have.

OpenStudy (anonymous):

thankyou! I've had so much trouble with that question!

OpenStudy (mertsj):

Something is very wrong. Are you sure you posted the question EXACTLY as stated?

OpenStudy (anonymous):

yes! I can post a screenshot if you'd like. I just don't know how to upload a picture

OpenStudy (mertsj):

That would be good.

OpenStudy (anonymous):

do you know how I can upload a pic?

OpenStudy (mertsj):

All circles have x^2 and y^2 in their equations.

OpenStudy (anonymous):

yeah that's what I thought but this question is really weird and theres a lot similiar to these on my test

OpenStudy (mertsj):

I have never done that on this site.

OpenStudy (anonymous):

I've seen other people do it...hmmm let me upload the pic to a site and then share the link

OpenStudy (mertsj):

ok

OpenStudy (mertsj):

It hasn't finished loading yet. I should warn you that my internet often goes down at 11 for some strange reason the Frontier can't seem to figure out. But the way the problem is posted, it has to be either choice b or .

OpenStudy (anonymous):

Oh aw alright! What time is it rn where ur at? I chose answer b but I haven't submitted it yet

OpenStudy (mertsj):

I see what you mean. For the second problem, let's try choice b. 10:52

OpenStudy (mathstudent55):

The equation of a circle must have both the x-term and the y-term squared. The equation you were given is not a circle. I think they left out the y squared term.

OpenStudy (anonymous):

aw man thats soon :( oh well...... wait for the second problem? so not the first one right? @mathstudent55 there are many questions like this on my test :(

OpenStudy (mertsj):

\[(x+4)^2+(y-3)^2=(\sqrt{34})^2\] \[x^2+8x+16+y^2-6y+9=34\] \[x^2+8x+y^2-6y=9\]

OpenStudy (anonymous):

so b for that one?

OpenStudy (mertsj):

I would choose b and then I would point out to my teacher that these equations cannot be circles.

OpenStudy (anonymous):

Oh okay! Thankyou so much for your time and help!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

do you know how to award a medal ? I'm not really sure how to haha

OpenStudy (anonymous):

nvm I think i got it